# Well-posedness of stochastic mSQG equations with Kraichnan noise and $L^{p}$ data

### v1

@@ 3.2 A priori energy estimate
$\begin{split}&\langle\theta_{t}^{\delta},G_{\beta}^{\delta}*\theta_{t}^{\delta}\rangle-\int_{0}^{t}\iint\limits_{\mathbb{R}^{2}\times\mathbb{R}^{2}}\mathrm{tr}\big{[}(Q^{\delta}(0)-Q^{\delta}(x-y))D^{2}G_{\beta}^{\delta}(x-y)\big{]}\theta_{r}^{\delta}(x)\theta_{r}^{\delta}(y)\,\mathrm{d}x\mathrm{d}y\mathrm{d}r\\ &=\langle\theta_{0}^{\delta},G_{\beta}^{\delta}*\theta_{0}^{\delta}\rangle-2\int_{0}^{t}\sum_{k}\langle\sigma_{k}^{\delta}\cdot\nabla\theta_{r}^{\delta},G_{\beta}^{\delta}*\theta_{r}^{\delta}\rangle\,\mathrm{d}B_{r}^{k}.\end{split}$ (3.2)
@@
$\begin{split}&\mathbb{E}[\langle\theta_{t}^{\delta},G_{\beta}^{\delta}*\theta_{t}^{\delta}\rangle]-\langle\theta_{0}^{\delta},G_{\beta}^{\delta}*\theta_{0}^{\delta}\rangle\\ =&\int_{0}^{t}\mathbb{E}\iint\limits_{\mathbb{R}^{2}\times\mathbb{R}^{2}}\mathrm{tr}\big{[}(Q^{\delta}(0)-Q^{\delta}(x-y))D^{2}G_{\beta}^{\delta}(x-y)\big{]}\theta_{r}^{\delta}(x)\theta_{r}^{\delta}(y)\,\mathrm{d}x\mathrm{d}y\mathrm{d}r.\end{split}$ (3.3)
@@
Consider the functional $F:H^{-4}\ni\theta\rightarrow\langle\theta,G_{\beta}^{\delta}*\theta\rangle\in\mathbb{R}$, whose Fréchet derivatives are
$DF(\theta)v$ $=2\langle v,G_{\beta}^{\delta}*\theta\rangle,$
$D^{2}F(\theta)[v,w]$ $=2\langle v,G_{\beta}^{\delta}*w\rangle.$
The classical Itô formula on Hilbert space (see [13, Theorem 4.17]) implies that
$\begin{split}\mathrm{d}\langle\theta^{\delta},G_{\beta}^{\delta}*\theta^{\delta}\rangle&=-2\big{\langle}(K_{\beta}^{\delta}*\theta^{\delta})\cdot\nabla\theta^{\delta},G_{\beta}^{\delta}*\theta^{\delta}\big{\rangle}\,\mathrm{d}t-2\sum_{k}\langle\sigma_{k}^{\delta}\cdot\nabla\theta^{\delta},G_{\beta}^{\delta}*\theta^{\delta}\rangle\,\mathrm{d}B^{k}\\ &\quad+\langle c_{\delta}\Delta\theta^{\delta},G_{\beta}^{\delta}*\theta^{\delta}\rangle\,\mathrm{d}t+\sum_{k}\langle\sigma_{k}^{\delta}\cdot\nabla\theta^{\delta},G_{\beta}^{\delta}*(\sigma_{k}^{\delta}\cdot\nabla\theta^{\delta})\rangle\,\mathrm{d}t.\end{split}$ (3.4)
@@
$\langle(K_{\beta}^{\delta}*\theta^{\delta})\cdot\nabla\theta^{\delta},G_{\beta}^{\delta}*\theta^{\delta}\rangle=-\langle\theta^{\delta},\nabla^{\perp}(G_{\beta}^{\delta}*\theta^{\delta})\cdot\nabla(G_{\beta}^{\delta}*\theta^{\delta})\rangle=0.$
@@
$\mathbb{E}\sum_{k}|\langle\sigma_{k}^{\delta}\cdot\nabla\theta^{\delta},G_{\beta}^{\delta}*\theta^{\delta}\rangle|^{2}$ $=\mathbb{E}\iint\limits_{\mathbb{R}^{2}\times\mathbb{R}^{2}}\theta^{\delta}(x)\theta^{\delta}(y)\nabla G_{\beta}^{\delta}*\theta^{\delta}(x)\cdot Q^{\delta}(x-y)\nabla G_{\beta}^{\delta}*\theta^{\delta}(y)\,\mathrm{d}x\mathrm{d}y$
@@
$\langle c_{\delta}\Delta\theta^{\delta},G_{\beta}^{\delta}*\theta^{\delta}\rangle=\langle\theta^{\delta},c_{\delta}\Delta G_{\beta}^{\delta}*\theta^{\delta}\rangle=\iint\limits_{\mathbb{R}^{2}\times\mathbb{R}^{2}}\mathrm{tr}\big{[}Q^{\delta}(0)D^{2}G_{\beta}^{\delta}(x-y)\big{]}\theta^{\delta}(x)\theta^{\delta}(y)\,\mathrm{d}x\mathrm{d}y,$
@@
$\sum_{k}\langle\sigma_{k}^{\delta}\cdot\nabla\theta^{\delta},G_{\beta}^{\delta}*(\sigma_{k}^{\delta}\cdot\nabla\theta^{\delta})\rangle$ $=\sum_{k}\sum_{i,j}\langle\partial_{i}(\sigma_{k}^{\delta,i}\theta^{\delta}),\partial_{j}G_{\beta}^{\delta}*(\sigma_{k}^{\delta,j}\theta^{\delta})\rangle$
$=-\sum_{i,j}\sum_{k}\langle\sigma_{k}^{\delta,i}\theta^{\delta},\partial_{ij}^{2}G_{\beta}^{\delta}*(\sigma_{k}^{\delta,j}\theta^{\delta})\rangle$
@@
Suppose $\alpha+\beta>2$, then for the term $A(x)=\mathrm{tr}\big{[}(Q(0)-Q(x))D^{2}G_{\beta}(x)\big{]}\varphi(x)$, there exist two positive constant $c,C$ such that
@@
Recalling the fact $\beta_{N}=(1+\alpha)\beta_{L}$ in Lemma 2.2 and the definition of Riesz kernel, we have
$\begin{split}\mathrm{tr}\big{[}(Q(0)-Q(x))D^{2}G_{\beta}(x)\big{]}=&-\frac{2-\beta}{\gamma(\beta)}(\alpha+\beta-2)\beta_{L}\gamma(2\alpha+\beta-2)G_{2\alpha+\beta-2}(x)\\ \quad&+\frac{2-\beta}{\gamma(\beta)}\frac{1}{|x|^{4-\beta}}\big{(}(3-\beta)\mathrm{Rem}_{1-u^{2}}(|x|)-\mathrm{Rem}_{u^{2}}(|x|)\big{)}.\end{split}$ (3.6)
For the control of $\varphi G_{2\alpha+\beta-2}$ in Fourier modes, by [10, Lemma 4.3], for some constant $C$, we have
@@
Now choosing $\bar{\delta}=(\alpha+\beta-2)\frac{(2-\beta)\gamma(2\alpha+\beta-2)}{4(2\pi)^{2\alpha+\beta-2}\gamma(\beta)}\beta_{L}$, we conclude that for some constant $C>0$,
$\hat{A}(\xi)\leq-(\alpha+\beta-2)\frac{(2-\beta)\gamma(2\alpha+\beta-2)\beta_{L}}{4(2\pi)^{2\alpha+\beta-2}\gamma(\beta)}\langle\xi\rangle^{2-2\alpha-\beta}+C\langle\xi\rangle^{-\beta}.$
@@
Let $\alpha\in(0,1)$ and $\beta\in(1,2)$ satisfy $\alpha+\beta>2$. There exist constants $c,C>0$ such that $\mathbb{P}$-a.s., for every $t\in[0,T]$,
@@
Let $\alpha\in(0,1)$ and $\beta\in(1,2)$ satisfy $\alpha+\beta>2$, then there exist constants $c,C>0$ such that, for every $10<\delta<$ 1,
@@ 3.3 Bounds on time continuity
then by the property of Riesz transform and Hardy-Littlewood-Sobolev inequality (see e.g. [39, Chapter III, Section 1] and [3, Theorem 1.8]), we have $u^{\delta}$ is in $L^{p_{*}}$ and $\|u^{\delta}\|_{L^{p_{*}}}\lesssim\|\theta^{\delta}\|_{L^{p}}$, where $p_{*}^{-1}=p^{-1}-\frac{\beta-1}{2}$. Now let $q^{-1}=p^{-1}+p_{*}^{-1}$, by the condition $\frac{4}{\beta+1}\leq p\leq 2$, we know $21\leq q<$ 2. Sobolev embedding implies that $L^{q}\hookrightarrow H^{-2}$, hence together with Hölder inequality, we obtain
$\mathbb{E}\bigg{[}\sup\limits_{r\in[0,T]}\|u_{r}^{\delta}\theta_{r}^{\delta}\|_{H^{-2}}^{\lambda}\bigg{]}$ $\lesssim\mathbb{E}\bigg{[}\sup\limits_{r\in[0,T]}\|u_{r}^{\delta}\theta_{r}^{\delta}\|_{L^{q}}^{\lambda}\bigg{]}\lesssim\mathbb{E}\bigg{[}\sup\limits_{r\in[0,T]}\|u_{r}^{\delta}\|_{L^{p_{*}}}^{\lambda}\|\theta_{r}^{\delta}\|_{L^{p}}^{\lambda}\bigg{]}$
@@
$\begin{split}\mathbb{E}\bigg{[}\sup\limits_{r\in[0,T]}\|(u_{r}^{\delta}-\tilde{u}_{r}^{\delta})\theta_{r}^{\delta}\|_{L^{2}}^{\lambda}\bigg{]}&\leq\mathbb{E}\bigg{[}\sup\limits_{r\in[0,T]}\|u_{r}^{\delta}-\tilde{u}_{r}^{\delta}\|_{L^{\infty}}^{\lambda}\|\theta_{r}^{\delta}\|_{L^{2}}^{\lambda}\bigg{]}\\ &\leq\|\theta_{0}^{\delta}\|_{L^{2}}^{\lambda}\mathbb{E}\bigg{[}\sup\limits_{r\in[0,T]}\|(\nabla G_{\beta}^{\delta}-\nabla G_{\beta})*\theta_{r}^{\delta}\|_{L^{\infty}}^{\lambda}\bigg{]}\\ &\lesssim\Big{[}\delta^{\frac{\beta-1}{\beta+3}}\|\theta_{0}^{\delta}\|_{L^{1}}^{1/2}\|\theta_{0}^{\delta}\|_{L^{\infty}}^{1/2}(\|\theta_{0}^{\delta}\|_{L^{\infty}}+\|\theta_{0}^{\delta}\|_{L^{1}})\Big{]}^{\lambda}\\ &\lesssim\Big{[}\delta^{\frac{\beta-1}{\beta+3}}(\|\theta_{0}^{\delta}\|_{L^{\infty}}+\|\theta_{0}^{\delta}\|_{L^{1}})^{2}\Big{]}^{\lambda}=o(1).\end{split}$ (3.9)
@@
$\sum_{k}\|\sigma_{k}^{\delta}\theta^{\delta}\|_{H^{-3}}^{2}=\int\theta^{\delta}(x)(\psi*\theta^{\delta})(x)\,\mathrm{d}x\\ =\int\big{|}\widehat{\theta^{\delta}}(\xi)\big{|}^{2}\hat{\psi}(\xi)\,\mathrm{d}\xi.$ (3.11)
@@
By the Kolmogorov criterion (see e.g. [32, Theorem 2.9]), we conclude that, for every $20<\gamma<1/$ 2,
@@ 4 Weak existence
The key point of the argument is the uniform bound on the $H^{\frac{\beta}{2}-\alpha}$ norm of the velocity field $u^{\delta}=K_{\beta}*\theta^{\delta}$. Indeed, for $\beta>2\alpha$, the embedding $H^{\frac{\beta}{2}-\alpha}\hookrightarrow L^{2}$ is compact on every bounded domain of $\mathbb{R}^{2}$ and hence the uniform $H^{\frac{\beta}{2}-\alpha}$ bound implies convergence in strong $L_{loc}^{2}$ topology, which allows to pass to the limit in the nonlinear term.
@@
The following lemma is from [10, Lemma 6.1].
@@
In fact, for every $\epsilon>0$, the space $H^{\frac{\beta}{2}-\alpha}(\mathbb{R}^{2})$ compactly embedds into $H_{loc}^{\frac{\beta}{2}-\alpha-\epsilon}(\mathbb{R}^{2})$. For details, see e.g. [21, Lemma A.4].
@@
Let $u^{\delta}=K_{\beta}*\theta^{\delta}$. We will prove the tightness of the laws of the family $(u^{\delta})_{\delta}$ in the space $L_{t}^{2}(L^{2}(\mathbb{R}^{2};w))$. Recall the following facts:
$\|K_{\beta}*f\|_{\dot{H}^{-1+\frac{\beta}{2}}}=(2\pi)^{1-\beta}\|f\|_{\dot{H}^{-\frac{\beta}{2}}},$
$\|K_{\beta}*f\|_{H^{-5+\beta}}=(2\pi)^{1-\beta}\|f\|_{\tilde{H}^{-4}}.$
@@
$\begin{split}\|K_{\beta}*f\|_{\dot{H}^{\frac{\beta}{2}-\alpha}}^{2}&=(2\pi)^{2-2\beta}\int_{\mathbb{R}^{2}}|\xi|^{-\beta+2-2\alpha}|\hat{f}(\xi)|^{2}\,\mathrm{d}\xi\\ &\lesssim\int_{|\xi|\leq 1}|\xi|^{-\beta}|\hat{f}(\xi)|^{2}\,\mathrm{d}\xi+\int_{|\xi|\geq 1}\langle\xi\rangle^{-\beta+2-2\alpha}|\hat{f}(\xi)|^{2}\mathrm{d}\xi\\ &\lesssim\|f\|_{\dot{H}^{-\frac{\beta}{2}}}^{2}+\|f\|_{H^{-\frac{\beta}{2}+1-\alpha}}^{2}.\end{split}$
@@
$\|K_{\beta}*f\|_{L^{2}}\leq\|K_{\beta}*f\|_{\dot{H}^{-1+\frac{\beta}{2}}}^{1-\lambda}\|K_{\beta}*f\|_{\dot{H}^{\frac{\beta}{2}-\alpha}}^{\lambda}\leq\|K_{\beta}*f\|_{\dot{H}^{-1+\frac{\beta}{2}}}+\|K_{\beta}*f\|_{\dot{H}^{\frac{\beta}{2}-\alpha}},$
@@
$\|K_{\beta}*f\|_{H^{\frac{\beta}{2}-\alpha}}^{2}=\|K_{\beta}*f\|_{L^{2}}^{2}+\|K_{\beta}*f\|_{\dot{H}^{\frac{\beta}{2}-\alpha}}^{2}\lesssim\|f\|_{\dot{H}^{-\frac{\beta}{2}}}^{2}+\|f\|_{H^{-\frac{\beta}{2}+1-\alpha}}^{2}.$
@@
By Skorohod representation theorem [27, Theorem 5.31], there exists a complete probability space $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{P}})$, a sequence of $L_{t}^{2}(L^{2}(\mathbb{R}^{2});w)\times C_{t}^{\mathbb{N}}$-valued random variables $(\tilde{u}^{\delta_{n}},(\tilde{B}^{k,\delta_{n}})_{k})$ on $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{P}})$, with $\delta_{n}\rightarrow 0$ as $n\rightarrow\infty$, and a $L_{t}^{2}(L^{2}(\mathbb{R}^{2});w)\times C_{t}^{\mathbb{N}}$-valued random variable $(\tilde{u},(\tilde{B}^{k})_{k})$ such that each $(\tilde{u}^{\delta_{n}},(\tilde{B}^{k,\delta_{n}})_{k})$ has the same law as $(u^{\delta_{n}},(B^{k})_{k})$ and the sequence $(\tilde{u}^{\delta_{n}},(\tilde{B}^{k,\delta_{n}})_{k})_{n}$ converges $\tilde{\mathbb{P}}$-a.s. in the topology of $L_{t}^{2}(L^{2}(\mathbb{R}^{2});w)\times C_{t}^{\mathbb{N}}$ to $(\tilde{u},(\tilde{B}^{k})_{k})$ as $n\rightarrow\infty$. We call

### v2

@@ 3.2 A priori energy estimate
$\begin{split}&\langle\theta_{t}^{\delta},G_{\beta}^{\delta}\ast\theta_{t}^{\delta}\rangle-\int_{0}^{t}\iint\limits_{\mathbb{R}^{2}\times\mathbb{R}^{2}}\mathrm{tr}\big{[}(Q^{\delta}(0)-Q^{\delta}(x-y))D^{2}G_{\beta}^{\delta}(x-y)\big{]}\theta_{r}^{\delta}(x)\theta_{r}^{\delta}(y)\,\mathrm{d}x\mathrm{d}y\mathrm{d}r\\ &=\langle\theta_{0}^{\delta},G_{\beta}^{\delta}\ast\theta_{0}^{\delta}\rangle-2\int_{0}^{t}\sum_{k}\langle\sigma_{k}^{\delta}\cdot\nabla\theta_{r}^{\delta},G_{\beta}^{\delta}\ast\theta_{r}^{\delta}\rangle\,\mathrm{d}B_{r}^{k}.\end{split}$ (3.2)
@@
$\begin{split}&\mathbb{E}[\langle\theta_{t}^{\delta},G_{\beta}^{\delta}\ast\theta_{t}^{\delta}\rangle]-\langle\theta_{0}^{\delta},G_{\beta}^{\delta}\ast\theta_{0}^{\delta}\rangle\\ &=\int_{0}^{t}\mathbb{E}\iint\limits_{\mathbb{R}^{2}\times\mathbb{R}^{2}}\mathrm{tr}\big{[}(Q^{\delta}(0)-Q^{\delta}(x-y))D^{2}G_{\beta}^{\delta}(x-y)\big{]}\theta_{r}^{\delta}(x)\theta_{r}^{\delta}(y)\,\mathrm{d}x\mathrm{d}y\mathrm{d}r.\end{split}$ (3.3)
@@
Consider the functional $F:H^{-4}\ni\theta\rightarrow\langle\theta,G_{\beta}^{\delta}\ast\theta\rangle\in\mathbb{R}$, whose Fréchet derivatives are
$DF(\theta)v$ $=2\langle v,G_{\beta}^{\delta}\ast\theta\rangle,$
$D^{2}F(\theta)[v,w]$ $=2\langle v,G_{\beta}^{\delta}\ast w\rangle.$
The classical Itô formula on Hilbert space (see [14, Theorem 4.17]) implies that
$\begin{split}\mathrm{d}\langle\theta^{\delta},G_{\beta}^{\delta}\ast\theta^{\delta}\rangle&=-2\big{\langle}(K_{\beta}^{\delta}\ast\theta^{\delta})\cdot\nabla\theta^{\delta},G_{\beta}^{\delta}\ast\theta^{\delta}\big{\rangle}\,\mathrm{d}t-2\sum_{k}\langle\sigma_{k}^{\delta}\cdot\nabla\theta^{\delta},G_{\beta}^{\delta}\ast\theta^{\delta}\rangle\,\mathrm{d}B^{k}\\ &\quad+\langle c_{\delta}\Delta\theta^{\delta},G_{\beta}^{\delta}\ast\theta^{\delta}\rangle\,\mathrm{d}t+\sum_{k}\langle\sigma_{k}^{\delta}\cdot\nabla\theta^{\delta},G_{\beta}^{\delta}\ast(\sigma_{k}^{\delta}\cdot\nabla\theta^{\delta})\rangle\,\mathrm{d}t.\end{split}$ (3.4)
@@
$\langle(K_{\beta}^{\delta}\ast\theta^{\delta})\cdot\nabla\theta^{\delta},G_{\beta}^{\delta}\ast\theta^{\delta}\rangle=-\langle\theta^{\delta},\nabla^{\perp}(G_{\beta}^{\delta}\ast\theta^{\delta})\cdot\nabla(G_{\beta}^{\delta}\ast\theta^{\delta})\rangle=0.$
@@
$\mathbb{E}\sum_{k}|\langle\sigma_{k}^{\delta}\cdot\nabla\theta^{\delta},G_{\beta}^{\delta}\ast\theta^{\delta}\rangle|^{2}$ $=\mathbb{E}\iint\limits_{\mathbb{R}^{2}\times\mathbb{R}^{2}}\theta^{\delta}(x)\theta^{\delta}(y)\nabla G_{\beta}^{\delta}\ast\theta^{\delta}(x)\cdot Q^{\delta}(x-y)\nabla G_{\beta}^{\delta}\ast\theta^{\delta}(y)\,\mathrm{d}x\mathrm{d}y$
@@
$\langle c_{\delta}\Delta\theta^{\delta},G_{\beta}^{\delta}\ast\theta^{\delta}\rangle=\langle\theta^{\delta},c_{\delta}\Delta G_{\beta}^{\delta}\ast\theta^{\delta}\rangle=\iint\limits_{\mathbb{R}^{2}\times\mathbb{R}^{2}}\mathrm{tr}\big{[}Q^{\delta}(0)D^{2}G_{\beta}^{\delta}(x-y)\big{]}\theta^{\delta}(x)\theta^{\delta}(y)\,\mathrm{d}x\mathrm{d}y,$
@@
$\sum_{k}\langle\sigma_{k}^{\delta}\cdot\nabla\theta^{\delta},G_{\beta}^{\delta}\ast(\sigma_{k}^{\delta}\cdot\nabla\theta^{\delta})\rangle$ $=\sum_{k}\sum_{i,j}\langle\partial_{i}(\sigma_{k}^{\delta,i}\theta^{\delta}),\partial_{j}G_{\beta}^{\delta}\ast(\sigma_{k}^{\delta,j}\theta^{\delta})\rangle$
$=-\sum_{i,j}\sum_{k}\langle\sigma_{k}^{\delta,i}\theta^{\delta},\partial_{ij}^{2}G_{\beta}^{\delta}\ast(\sigma_{k}^{\delta,j}\theta^{\delta})\rangle$
@@
Suppose $22\alpha+\beta>$ 2, then for the term $A(x)=\mathrm{tr}\big{[}(Q(0)-Q(x))D^{2}G_{\beta}(x)\big{]}\varphi(x)$, there exist two positive constants $c,C$ such that
@@
Recalling the fact $\beta_{N}=(1+2\alpha)\beta_{L}$ in Lemma 2.2 and the definition of Riesz kernel, we have
$\begin{split}\mathrm{tr}\big{[}(Q(0)-Q(x))D^{2}G_{\beta}(x)\big{]}=&-\frac{2-\beta}{\gamma(\beta)}(2\alpha+\beta-2)\beta_{L}\gamma(2\alpha+\beta-2)G_{2\alpha+\beta-2}(x)\\ \quad&+\frac{2-\beta}{\gamma(\beta)}\frac{1}{|x|^{4-\beta}}\big{(}(3-\beta)\mathrm{Rem}_{1-u^{2}}(|x|)-\mathrm{Rem}_{u^{2}}(|x|)\big{)}.\end{split}$ (3.6)
For the control of $\varphi G_{2\alpha+\beta-2}$ in Fourier modes, by [11, Lemma 4.3], there exists some constant $C>0$ such that
@@
Now choosing $\bar{\delta}=(2\alpha+\beta-2)\frac{(2-\beta)\gamma(2\alpha+\beta-2)}{4(2\pi)^{2\alpha+\beta-2}\gamma(\beta)}\beta_{L}$, we conclude that for some constant $C>0$,
$\hat{A}(\xi)\leq-(2\alpha+\beta-2)\frac{(2-\beta)\gamma(2\alpha+\beta-2)\beta_{L}}{4(2\pi)^{2\alpha+\beta-2}\gamma(\beta)}\langle\xi\rangle^{2-2\alpha-\beta}+C\langle\xi\rangle^{-\beta}.$
@@
Let $\alpha\in(0,1)$ and $\beta\in(1,2)$ satisfy $22\alpha+\beta>$ 2. There exist constants $c,C>0$ such that $\mathbb{P}$-a.s., for every $t\in[0,T]$,
@@
Let $\alpha\in(0,1)$ and $\beta\in(1,2)$ satisfy $22\alpha+\beta>$ 2, then there exist constants $c,\,C>0$ such that, for every $10<\delta<$ 1,
@@ 3.3 Bounds on time continuity
then by the property of Riesz transform and Hardy-Littlewood-Sobolev inequality (see e.g. [40, Chapter III, Section 1] and [4, Theorem 1.8]), we have $u^{\delta}$ is in $L^{p_{\ast}}$ and $\|u^{\delta}\|_{L^{p_{\ast}}}\lesssim\|\theta^{\delta}\|_{L^{p}}$, where $p_{\ast}^{-1}=p^{-1}-\frac{\beta-1}{2}$. Now let $q^{-1}=p^{-1}+p_{\ast}^{-1}$, by the condition $\frac{4}{\beta+1}\leq p\leq 2$, we know $21\leq q<$ 2. Sobolev embedding implies that $L^{q}\hookrightarrow H^{-2}$, hence together with Hölder inequality, we obtain
$\mathbb{E}\bigg{[}\sup\limits_{r\in[0,T]}\|u_{r}^{\delta}\theta_{r}^{\delta}\|_{H^{-2}}^{\lambda}\bigg{]}$ $\lesssim\mathbb{E}\bigg{[}\sup\limits_{r\in[0,T]}\|u_{r}^{\delta}\theta_{r}^{\delta}\|_{L^{q}}^{\lambda}\bigg{]}\lesssim\mathbb{E}\bigg{[}\sup\limits_{r\in[0,T]}\|u_{r}^{\delta}\|_{L^{p_{\ast}}}^{\lambda}\|\theta_{r}^{\delta}\|_{L^{p}}^{\lambda}\bigg{]}$
@@
$\begin{split}\mathbb{E}\bigg{[}\sup\limits_{r\in[0,T]}\|(u_{r}^{\delta}-\tilde{u}_{r}^{\delta})\theta_{r}^{\delta}\|_{L^{2}}^{\lambda}\bigg{]}&\leq\mathbb{E}\bigg{[}\sup\limits_{r\in[0,T]}\|u_{r}^{\delta}-\tilde{u}_{r}^{\delta}\|_{L^{\infty}}^{\lambda}\|\theta_{r}^{\delta}\|_{L^{2}}^{\lambda}\bigg{]}\\ &\leq\|\theta_{0}^{\delta}\|_{L^{2}}^{\lambda}\mathbb{E}\bigg{[}\sup\limits_{r\in[0,T]}\|(\nabla G_{\beta}^{\delta}-\nabla G_{\beta})\ast\theta_{r}^{\delta}\|_{L^{\infty}}^{\lambda}\bigg{]}\\ &\lesssim\Big{[}\delta^{\frac{\beta-1}{\beta+3}}\|\theta_{0}^{\delta}\|_{L^{1}}^{1/2}\|\theta_{0}^{\delta}\|_{L^{\infty}}^{1/2}(\|\theta_{0}^{\delta}\|_{L^{\infty}}+\|\theta_{0}^{\delta}\|_{L^{1}})\Big{]}^{\lambda}\\ &\lesssim\Big{[}\delta^{\frac{\beta-1}{\beta+3}}(\|\theta_{0}^{\delta}\|_{L^{\infty}}+\|\theta_{0}^{\delta}\|_{L^{1}})^{2}\Big{]}^{\lambda}=o(1).\end{split}$ (3.9)
@@
$\sum_{k}\|\sigma_{k}^{\delta}\theta^{\delta}\|_{H^{-3}}^{2}=\int\theta^{\delta}(x)(\psi\ast\theta^{\delta})(x)\,\mathrm{d}x\\ =\int\big{|}\widehat{\theta^{\delta}}(\xi)\big{|}^{2}\hat{\psi}(\xi)\,\mathrm{d}\xi.$ (3.11)
@@
By the Kolmogorov criterion (see e.g. [33, Theorem 2.9]), we conclude that, for every $20<\gamma<1/$ 2,
@@ 4 Weak existence
The key point of the argument is the uniform bound on the $H^{\frac{\beta}{2}-\alpha}$ norm of the velocity field $u^{\delta}=K_{\beta}\ast\theta^{\delta}$. Indeed, for $\beta>2\alpha$, the embedding $H^{\frac{\beta}{2}-\alpha}\hookrightarrow L^{2}$ is compact on every bounded domain of $\mathbb{R}^{2}$ and hence the uniform $H^{\frac{\beta}{2}-\alpha}$ bound implies convergence in strong $L_{loc}^{2}$ topology, which allows to pass to the limit in the nonlinear term.
@@
The following lemma is from [11, Lemma 6.1].
@@
In fact, for every $\epsilon>0$, the space $H^{\frac{\beta}{2}-\alpha}(\mathbb{R}^{2})$ compactly embedds into $H_{loc}^{\frac{\beta}{2}-\alpha-\epsilon}(\mathbb{R}^{2})$. For details, see e.g. [22, Lemma A.4].
@@
Let $u^{\delta}=K_{\beta}\ast\theta^{\delta}$. We will prove the tightness of the laws of the family $(u^{\delta})_{\delta}$ in the space $L_{t}^{2}(L^{2}(\mathbb{R}^{2};w))$. Recall the following facts:
$\|K_{\beta}\ast f\|_{\dot{H}^{-1+\frac{\beta}{2}}}=(2\pi)^{1-\beta}\|f\|_{\dot{H}^{-\frac{\beta}{2}}},$
$\|K_{\beta}\ast f\|_{H^{-5+\beta}}=(2\pi)^{1-\beta}\|f\|_{\tilde{H}^{-4}}.$
@@
$\begin{split}\|K_{\beta}\ast f\|_{\dot{H}^{\frac{\beta}{2}-\alpha}}^{2}&=(2\pi)^{2-2\beta}\int_{\mathbb{R}^{2}}|\xi|^{-\beta+2-2\alpha}|\hat{f}(\xi)|^{2}\,\mathrm{d}\xi\\ &\lesssim\int_{|\xi|\leq 1}|\xi|^{-\beta}|\hat{f}(\xi)|^{2}\,\mathrm{d}\xi+\int_{|\xi|\geq 1}\langle\xi\rangle^{-\beta+2-2\alpha}|\hat{f}(\xi)|^{2}\mathrm{d}\xi\\ &\lesssim\|f\|_{\dot{H}^{-\frac{\beta}{2}}}^{2}+\|f\|_{H^{-\frac{\beta}{2}+1-\alpha}}^{2}.\end{split}$
@@
$\|K_{\beta}\ast f\|_{L^{2}}\leq\|K_{\beta}\ast f\|_{\dot{H}^{-1+\frac{\beta}{2}}}^{1-\lambda}\|K_{\beta}\ast f\|_{\dot{H}^{\frac{\beta}{2}-\alpha}}^{\lambda}\leq\|K_{\beta}\ast f\|_{\dot{H}^{-1+\frac{\beta}{2}}}+\|K_{\beta}\ast f\|_{\dot{H}^{\frac{\beta}{2}-\alpha}},$
@@
$\|K_{\beta}\ast f\|_{H^{\frac{\beta}{2}-\alpha}}^{2}=\|K_{\beta}\ast f\|_{L^{2}}^{2}+\|K_{\beta}\ast f\|_{\dot{H}^{\frac{\beta}{2}-\alpha}}^{2}\lesssim\|f\|_{\dot{H}^{-\frac{\beta}{2}}}^{2}+\|f\|_{H^{-\frac{\beta}{2}+1-\alpha}}^{2}.$
@@
By Skorohod representation theorem [28, Theorem 5.31], there exists a complete probability space $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{P}})$, a sequence of $L_{t}^{2}(L^{2}(\mathbb{R}^{2});w)\times C_{t}^{\mathbb{N}}$-valued random variables $(\tilde{u}^{\delta_{n}},(\tilde{B}^{k,\delta_{n}})_{k})$ on $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{P}})$, with $\delta_{n}\rightarrow 0$ as $n\rightarrow\infty$, and a $L_{t}^{2}(L^{2}(\mathbb{R}^{2});w)\times C_{t}^{\mathbb{N}}$-valued random variable $(\tilde{u},(\tilde{B}^{k})_{k})$ such that each $(\tilde{u}^{\delta_{n}},(\tilde{B}^{k,\delta_{n}})_{k})$ has the same law as $(u^{\delta_{n}},(B^{k})_{k})$ and the sequence $(\tilde{u}^{\delta_{n}},(\tilde{B}^{k,\delta_{n}})_{k})_{n}$ converges $\tilde{\mathbb{P}}$-a.s. in the topology of $L_{t}^{2}(L^{2}(\mathbb{R}^{2});w)\times C_{t}^{\mathbb{N}}$ to $(\tilde{u},(\tilde{B}^{k})_{k})$ as $n\rightarrow\infty$. We call